Change page style: 

NIRI Update: results from semester 2003A

You are here

19 Jun 2003

We recently completed the last scheduled NIRI observing block for semester 2003A. The weather cooperated and we consistently achieved 0.3 arcsec images using PWFS2. We were able to complete about 90% of the queue Band 1 observations.

A couple of months ago we started noticing intermittent failures of the pupil wheel again. Since we could no longer move it reliably, we were forced to move it into an open position and leave it there. The grisms are located in the pupil wheel, so locking it out effectively prohibited further spectroscopic observations. Unfortunately, we were unable to complete many of the highly-ranked spectroscopic observations.

To address this and other NIRI issues, we are beginning a NIRI shutdown of one month. NIRI will be opened in the instrument lab, the pupil wheel fixed, and several other mechanisms examined and tested. We will use the opportunity to install new grisms for adaptive optics spectroscopic observations with Altair, new slits, and new narrow band filters that have recently arrived. The web pages will be updated for these changes and additions soon. Please consult the web pages for updates when preparing Phase II proposals for 2003B.

The elevated read noise noticed earlier in the semester has been eliminated and the noise is back to the nominal levels expected. There is still some 60 Hz pattern noise, and we will be working to eliminate it over the summer.

NIRI was used in recent months to help commission Altair, the Gemini facility adaptive optics system. An example of the near-diffraction limited performace in the near-IR can be seen in the NIRI images taken with Altair.

Last update Jun. 19, 2003; Joe Jensen

Gemini Observatory Participants